Institut for Matematik og Datalogi Odense Universitet

{May 3, 1995 \atop JFB}

DM15 – F95 – Week 28

Lecture, May 1

We covered up through page 84 in the second chapter of Mitchell’s monograph.

Lecture, May 8

We will continue on the second chapter of Mitchell’s monograph, probably covering through section 2.5.1.

Lecture, May 15

We will continue on the second chapter of Mitchell’s monograph, covering up through section 2.5.5.

Problems to be discussed in week 20

• Suppose that M is a term in PCF. Show that if M =_{op} 1, then M =_{ax} 1.

• Find two PCF terms M and N, both of which are closed terms of observable type, without normal forms. M and N should satisfy the following: M =_{op} N, but M ≠_{ax} N.

• Do exercises 2.4.17 on page 87 of Mitchell. Note that the deterministic evaluator is an interpreter which uses the left-most reduction strategy.

• Do exercise 2.4.18 on page 87 of Mitchell.

• Do exercise 2.4.22 on page 91 of Mitchell.

• Do exercises 2.5.5 and 2.5.6 on page 95 of Mitchell. Do the necessary reductions to evaluate comp 2.

• Do exercise 2.5.10 on page 98 of Mitchell.

Announcement/Reminder

• There will be a meeting for faculty and students in computer science and applied math in the seminar room on Tuesday, May 9, at 14:15. The faculty will present the elective courses they will teach in the fall. The reformed program for the mandatory courses will also be explained, and Jørgen Bang-Jensen will talk about bachelor’s projects.